What are the steps to solve: (x-73)/x=.44
Answer by stanbon(75887)

What are the steps to solve: (x-73)/x=.44
x-73 = 0.44x
---
0.56x = 73
---
x = 73/0.56
---
x = 139.357....
Cheers,
Stan H.